# On the applicability of the cumulant expansion method for the calculation of transport properties in electron-phonon systems

### v2

@@ Abstract
We assess the accuracy of the cumulant expansion (CE) method, combined with the independent-particle approximation (IPA), for calculating charge mobility in electron–phonon systems. As representative testbeds, we consider the Peierls and Fröhlich models, which serve as simplified frameworks where accurate or numerically exact benchmarks are available. These are used to compare the CE results with those obtained using the Boltzmann formalism, the Migdal approximation, and its self consistent extension—approaches that are presently the most commonly employed alternatives for transport calculations. Supported by analytical arguments based on spectral sum rules and by our previous results for the Holstein model, we argue that, for weak to moderate coupling strengths and not-too-low temperatures, the CE within the IPA framework yields accurate results. In the case of the Peierls model, the role of vertex corrections is also discussed.
@@ I Introduction
First-principles calculations of charge carrier mobility in semiconductors with weak electron-phonon interaction have reached predictive power, mostly due to the advances in calculation of momentum-dependent electron-phonon coupling constants [1, 2]. In this regime the mobility calculation is typically performed within the Boltzmann semiclassical approach [3, 4, 5]. However, its underlying assumptions tend to break down at stronger couplings or higher temperatures, and the calculated mobility can thus strongly deviate from the experimental values [6]. In these cases, a significant improvement can be achieved by using the independent-particle approximation (IPA) within the Kubo formalism [7], combined with an appropriate method for calculating the single-particle Green’s function. One such choice is the Migdal approximation (MA), which retains only the lowest-order electron–phonon diagram, while higher-order effects can also be captured in the self-consistent version of this method (SCMA) [8].
@@
An effective strategy for such an assessment is to apply CE to well-controlled, simplified model systems [20, 21, 22], where accurate solutions are available and can serve as benchmarks [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. This is precisely the approach we adopted in our earlier studies, where we evaluated the accuracy of the CE method [37] and analyzed the role of vertex corrections [38] within the Holstein model [20]. This was made possible by the availability of several highly accurate approximate and entirely real-axis numerically exact methods for this model [39, 40, 41]. We found that, despite its simplicity, the CE method provided surprisingly accurate results for quasiparticle properties, spectral functions, and mobilities—not only at weak coupling, but also at intermediate strengths. Furthermore, in these regimes, vertex corrections were found to be relatively small [38]. However, it should be emphasized that the Holstein model assumes local (i.e., momentum-independent) electron-phonon coupling—a simplification that, while useful for gaining insight, limits the model’s physical realism. This highlights the importance of testing CE in more realistic models with momentum-dependent interactions.
@@ II.3 Kubo formula and independent particle approximation
II.4 Migdal and self-consistent Migdal approximation
@@
Figure 2: Self-energy in the Migdal approximation.
@@ III.2.1 One shot Migdal approximation
As we now show, the self-energy within the (one-shot) Migdal approximation (MA), given by Eq. (16), admits an analytic solution in the Peierls model. We begin with Eq. (16), perform the momentum shift $q\to q-k$, and note that only the even part of $|g_{k,q-k}|^{2}$ with respect to $q$ (see Eq. (29)) contributes. Consequently, the entire $q$-dependence in Eq. (16) comes from $\varepsilon_{q}$, so we can switch from momentum summation to energy integration
@@ III.4 HEOM method
The HEOM method provides a numerically exact density-matrix framework to study the dynamics of a system of interest [electrons in Eq. (1)] interacting with a collection of harmonic oscillators [phonons in Eq. (1)] [49, 50]. The method has enabled numerically exact computations of both single-electron [40] and transport [41, 38, 42, 43] properties in interacting electron–phonon models on a lattice in the limit of vanishing electron concentration. For completeness, here we give the main equations used to obtain HEOM results for the Peierls model (Sec. IV), while we refer the reader to the original papers for further details.
@@ III.4.2 Transport properties
To obtain the numerically exact mobility, it is convenient to recast the current autocorrelation function in Eq. (7) as $\langle j_{x}(t)j_{x}(0)\rangle=\mathrm{Tr}_{\mathrm{1e}}\left\{j_{x}\iota(t)\right\}$, where $\mathrm{Tr}_{\mathrm{1e}}$ denotes the trace over the subspace containing a single electron (and an arbitrary number of phonons) [38], while
$\iota(t)=\frac{(e^{-iHt}je^{-\beta H}e^{iHt})_{\mathrm{1e}}}{\mathrm{Tr}_{\mathrm{1e}}\>e^{-\beta H}}.$ (67)
The subscript “1e” in the numerator emphasizes that $\iota(t)$ is to be considered in the aforementioned subspace. The HEOM representation $\{\iota_{\mathbf{n}}^{(n)}(t)\}$ of $\iota(t)$ satisfies [42]
@@
In Eq. (68), $c_{\overline{m}}=n_{\mathrm{ph}}+\frac{1-(-1)^{m}}{2}$. At each instant $t\geq 0$, the current autocorrelation function reads [42]
$\begin{split}&\langle j_{x}(t)j_{x}(0)\rangle=\sum_{k}(\partial_{k}\varepsilon_{k})\langle k|\iota_{\mathbf{0}}^{(0)}(t)|k\rangle\\ &+\frac{1}{\sqrt{N}}\sum_{kqm}\sqrt{c_{m}}(\partial_{k}g_{k,q})\langle k|\iota_{\mathbf{0}_{qm}^{+}}^{(1)}(t)|k+q\rangle.\end{split}$ (69)
@@
$\begin{split}&\langle k|\iota_{\mathbf{n}}^{(n)}(0)|k+k_{\mathbf{n}}\rangle=(\partial_{k}\varepsilon_{k})\langle k|\rho_{\mathbf{n}}^{(n)}(0)|k+k_{\mathbf{n}}\rangle\\ &+\frac{1}{\sqrt{N}}\sum_{qm}\sqrt{1+n_{qm}}\sqrt{c_{m}}(\partial_{k}g_{k,-q}^{*})\langle k-q|\rho_{\mathbf{n}_{qm}^{+}}^{(n+1)}(0)|k+k_{\mathbf{n}}\rangle\\ &+\frac{1}{\sqrt{N}}\sum_{qm}\sqrt{n_{qm}}\sqrt{c_{m}}(\partial_{k}g_{k,q}^{*})\langle k+q|\rho_{\mathbf{n}_{qm}^{-}}^{(n-1)}(0)|k+k_{\mathbf{n}}\rangle.\end{split}$ (71)
In contrast to the HEOM for single-particle properties [Eq. (62)], the closing term in Eq. (68) is crucial in avoiding the long-time numerical instabilities characteristic for electron–phonon models with a finite number of phonon modes [41, 42, 52]. We always choose a sufficiently large $D$ so that the closing term merely stabilizes the long-time dynamics of the zeroth- and first-tier auxiliaries entering Eq. (69) without compromising the numerical exactness of the HEOM mobility [42]. We have also checked that using other closing terms [53, 51] does not change the mobility obtained by propagating Eq. (68[42, 43].
@@ IV.1.1 Quasiparticle properties
The results are shown in Fig. 4. Since most points in Fig. 4 lie in the higher-temperature range, where the mobility $\mu$ is small, we chose to plot $1/\mu(T)$ for better clarity. For completeness, the plot of $\mu(T)$ is provided in the following subsection; see Fig. 6($a_{2}$)–($c_{2}$). The curves labeled ”HEOM full” and ”HEOM IPA” represent numerically exact benchmarks for the mobility with [Eqs. (67)–(71)] and without [Eqs. (65) and (66)] vertex corrections, respectively. As seen from the figure, the discrepancy between the two clearly demonstrates that vertex corrections are not small, although the IPA result remains of the same order of magnitude as the full result. This difference is not solely due to the fact that $\mu_{x}$ vanishes within IPA [see Eq. (11)]; it is also evident in both the electronic [Panel (b)] and phononic [Panel (c)] contributions. In particular, while the IPA captures the qualitative temperature dependence of the electronic term in Panel (b), the phonon-assisted contribution in Panel (c) is not even qualitatively reproduced. We note that, for similar values of $\omega_{0},\lambda,$ and $T$, vertex corrections were found to be less pronounced in the Holstein model [38]. In contrast, their stronger impact in the Peierls model can be attributed to the pronounced momentum dependence of the electron–phonon matrix elements $g_{k,q}$ in both $k$ and $q$; see Eq. (25).
@@
All other methods (MA, SCMA, CE) are calculated within IPA, and thus need to be compared with HEOM IPA. As evident from the results, MA correctly captures the order of magnitude and the qualitative temperature dependence of $\mu$, but shows quantitative discrepancies when compared to HEOM IPA. In addition, the functional form of $\mu(T)$ at high temperatures, where the phonon-assisted contribution dominates, predicts a significantly steeper increase of $1/\mu$ with temperature than the benchmark. To improve upon the predictions of MA, one can use either SCMA or CE. Both of them are in very good agreement with HEOM IPA. However, while SCMA has self-consistency, CE is completely one-shot. This property makes it a computationally efficient and appealing alternative to other methods.
@@ IV.1.3 Mobility results in a wide range of parameter regimes
In Figs. 6 and 7, we present the mobility results over a wide range of parameter regimes. We do not discuss each panel individually, as the main conclusions are analogous to those from the previous subsection. Specifically: i) Vertex corrections are clearly visible but do not alter the order of magnitude of the results. In contrast to the Holstein model [38], however, we now also observe non-negligible vertex corrections even at weak interaction strengths [see, for example, Fig. 7($a_{1}$)]. This can be attributed to the strong dependence of the matrix elements $g_{k,q}$ on both $k$ and $q$. ii) The CE approach successfully reproduces the correct IPA result well beyond the range of validity of the Boltzmann quasiparticle picture. Unlike CE, which is a one-shot method, SCMA incorporates self-consistency, allowing it to achieve accurate results that also agree with the numerically exact HEOM IPA benchmark. Overall, both CE and SCMA provide a clear improvement over MA, which often deviates from the exact result except in the very weak coupling regime.
At lower temperatures, it becomes increasingly difficult to converge the CE results (this issue does not arise within SCMA). We verified that this problem is related to the long tail of the CE spectral functions extending toward negative frequencies, in close analogy with Fig. 5(c); see the corresponding spectral function in Sec. B of SM [44]. We also observed that this issue is even more pronounced at stronger coupling strengths. This behavior is expected, as the spectral functions become broader in this regime, thereby enhancing the contribution of the tail.
In these parameter regimes, the following issue arises for these reasons. Although it is sometimes possible, with considerable effort, to converge the CE mobility results with respect to the frequency cutoff $\Lambda$ at stronger couplings and lower temperatures, CE typically underestimates the HEOM IPA benchmark in these cases [see, for example, Fig. 6($a_{4}$)]. This can also be understood as a direct consequence of the spectral function tail. As discussed at the end of Sec. IV.1.2, the tail has the largest impact on $n_{e}$ (see Eq. (4)), making it unphysically large. Since $n_{e}$ appears in the denominator of the expressions for $\mu_{e}$ and $\mu_{ph}$ [see Eqs. (10) and (13)], these quantities within CE are systematically underestimated compared to the exact solution.

### v3

# Applicability of the cumulant expansion method for the calculation of transport properties in electron-phonon systems

@@ Abstract
We assess the accuracy of the cumulant expansion (CE) method, combined with the independent-particle approximation (IPA), for calculating charge mobility in electron–phonon systems. As representative testbeds, we consider the Peierls and Fröhlich models, which serve as simplified frameworks where accurate or numerically exact benchmarks are available. These are used to compare the CE results with those obtained using the Boltzmann formalism, the one-shot Migdal approximation, and its self consistent extension—approaches that are presently the most commonly employed alternatives for transport calculations. Supported by analytical arguments based on spectral sum rules and by our previous results for the Holstein model, we argue that, for weak to moderate coupling strengths and not-too-low temperatures, the CE within the IPA framework yields accurate results. In the case of the Peierls model, the role of vertex corrections is also discussed.
@@ I Introduction
First-principles calculations of charge carrier mobility in semiconductors with weak electron-phonon interaction have reached predictive power, mostly due to the advances in calculation of momentum-dependent electron-phonon coupling constants [1, 2]. In this regime the mobility calculation is typically performed within the Boltzmann semiclassical approach [3, 4, 5]. However, its underlying assumptions tend to break down at stronger couplings or higher temperatures, and the calculated mobility can thus strongly deviate from the experimental values [6]. The cases in which the Boltzmann approach is believed to break down include various materials where charge transport is governed by small polaron hopping [7, 8, 9, 10], perovskite oxides where charges must be treated beyond the quasiparticle picture [6, 11], and crystalline small-molecule organic semiconductors [12]. The limitations of the Boltzmann approach have been specifically demonstrated by recent calculations of spectral functions in real materials and the subsequent comparison of calculated mobilities to Boltzmann-derived results [13, 14]. In these cases, a significant improvement can be achieved by using the independent-particle approximation (IPA) within the Kubo formalism [15], combined with an appropriate method for calculating the single-particle Green’s function. One such choice is the one-shot Migdal approximation (MA), which retains only the lowest-order electron–phonon diagram, while higher-order effects can also be captured in the self-consistent version of this method (SCMA) [13].
@@
An effective strategy for such an assessment is to apply CE to well-controlled, simplified model systems [27, 28, 29], where accurate solutions are available and can serve as benchmarks [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. This is precisely the approach we adopted in our earlier studies, where we evaluated the accuracy of the CE method [45] and analyzed the role of vertex corrections [46] within the Holstein model [27]. This was made possible by the availability of several highly accurate approximate and entirely real-axis numerically exact methods for this model [47, 48, 49]. We found that, despite its simplicity, the CE method provided surprisingly accurate results for quasiparticle properties, spectral functions, and mobilities—not only at weak coupling, but also at intermediate strengths. Furthermore, in these regimes, vertex corrections were found to be relatively small [46]. However, it should be emphasized that the Holstein model assumes local (i.e., momentum-independent) electron-phonon coupling—a simplification that, while useful for gaining insight, limits the model’s physical realism. This highlights the importance of testing CE in more realistic models with momentum-dependent interactions.
@@ II.3 Kubo formula and independent particle approximation
II.4 One-shot Migdal and self-consistent Migdal approximation
@@
Figure 2: Self-energy in the one-shot Migdal approximation.
@@ III.2.1 One shot Migdal approximation
As we now show, the self-energy within the one-shot Migdal approximation (MA), given by Eq. (16), admits an analytic solution in the Peierls model. We begin with Eq. (16), perform the momentum shift $q\to q-k$, and note that only the even part of $|g_{k,q-k}|^{2}$ with respect to $q$ (see Eq. (29)) contributes. Consequently, the entire $q$-dependence in Eq. (16) comes from $\varepsilon_{q}$, so we can switch from momentum summation to energy integration
@@ III.4 HEOM method
The HEOM method provides a numerically exact density-matrix framework to study the dynamics of a system of interest [electrons in Eq. (1)] interacting with a collection of harmonic oscillators [phonons in Eq. (1)] [57, 58]. The method has enabled numerically exact computations of both single-electron [48] and transport [49, 46, 50, 51] properties in interacting electron–phonon models on a lattice in the limit of vanishing electron concentration. For completeness, here we give the actual equations that we have numerically implemented to obtain HEOM results for the Peierls model (Sec. IV). We refer the reader to the original papers for details on more formal aspects of the method, such as its formulation in the single-electron subspace [46, 47] or the generalized Wick’s theorem [50].
@@ III.4.2 Transport properties
To obtain the numerically exact mobility, it is convenient to recast the current autocorrelation function in Eq. (7) as $C_{jj}(t)=\langle j_{x}(t)j_{x}(0)\rangle/\widetilde{n}_{e}=\mathrm{Tr}_{\mathrm{1e}}\left\{j_{x}\iota(t)\right\}$, where $\mathrm{Tr}_{\mathrm{1e}}$ denotes the trace over the subspace containing a single electron (and an arbitrary number of phonons) [46], while
$\begin{split}\iota(t)&=\frac{(e^{-iHt}j_{x}e^{-\beta H}e^{iHt})_{\mathrm{1e}}}{\mathrm{Tr}_{\mathrm{1e}}\>e^{-\beta H}}\\ &=\iota_{e}(t)+\iota_{ph}(t).\end{split}$ (67)
In the first equality, the subscript “1e” in the numerator emphasizes that $\iota(t)$ is to be considered in the aforementioned subspace. The decomposition in the second equality follows from Eq. (8), where $\iota_{e}(t)$ $\iota_{ph}(t)$] is obtained by replacing $j_{x}$ in the first equality with its electronic (phonon-assisted) contribution. The HEOM representation $\{\iota_{\mathbf{n}}^{(n)}(t)\}$ of $\iota(t)$ satisfies [50]
@@
The very same equations govern the dynamics of the HEOM representations $\{\iota_{e,\mathbf{n}}^{(n)}(t)\}$ and $\{\iota_{ph,\mathbf{n}}^{(n)}(t)\}$ of $\iota_{e}(t)$ and $\iota_{ph}(t)$, respectively. In Eq. (68), $c_{\overline{m}}=n_{\mathrm{ph}}+\frac{1-(-1)^{m}}{2}$. At each instant $t\geq 0$, the current autocorrelation function reads [50]
$C_{jj}(t)=C_{jj,e}(t)+C_{jj,ph}(t)+C_{jj,x}(t),$ (69)
with the purely electronic contribution
$C_{jj,e}(t)=\sum_{k}(\partial_{k}\varepsilon_{k})\langle k|\iota_{e,\mathbf{0}}^{(0)}(t)|k\rangle,$ (70)
the phonon-assisted contribution
$C_{jj,ph}(t)=\frac{1}{\sqrt{N}}\sum_{kqm}\sqrt{c_{m}}(\partial_{k}g_{k,q})\langle k|\iota_{ph,\mathbf{0}_{qm}^{+}}^{(1)}(t)|k+q\rangle,$ (71)
and the cross contribution
$\begin{split}&C_{jj,x}(t)=\sum_{k}(\partial_{k}\varepsilon_{k})\langle k|\iota_{ph,\mathbf{0}}^{(0)}(t)|k\rangle\\ &+\frac{1}{\sqrt{N}}\sum_{kqm}\sqrt{c_{m}}(\partial_{k}g_{k,q})\langle k|\iota_{e,\mathbf{0}_{qm}^{+}}^{(1)}(t)|k+q\rangle.\end{split}$ (72)
Inserting Eqs. (70), (71), and (72) into Eq. (7), we obtain the HEOM full results $\mu_{e,\mathrm{HEOM}}^{\mathrm{full}}$, $\mu_{ph,\mathrm{HEOM}}^{\mathrm{full}}$, and $\mu_{x,\mathrm{HEOM}}^{\mathrm{full}}$ for the three contributions to the total mobility, see Eq. (9).
@@
$\begin{split}\langle k|\iota_{e,\mathbf{n}}^{(n)}(0)|k+k_{\mathbf{n}}\rangle=(\partial_{k}\varepsilon_{k})\langle k|\rho_{\mathbf{n}}^{(n,\mathrm{eq})}|k+k_{\mathbf{n}}\rangle,\end{split}$ (74)
$\begin{split}&\langle k|\iota_{ph,\mathbf{n}}^{(n)}(0)|k+k_{\mathbf{n}}\rangle=\\ &\frac{1}{\sqrt{N}}\sum_{qm}\sqrt{1+n_{qm}}\sqrt{c_{m}}(\partial_{k}g_{k,-q}^{*})\langle k-q|\rho_{\mathbf{n}_{qm}^{+}}^{(n+1,\mathrm{eq})}|k+k_{\mathbf{n}}\rangle\\ &+\frac{1}{\sqrt{N}}\sum_{qm}\sqrt{n_{qm}}\sqrt{c_{m}}(\partial_{k}g_{k,q}^{*})\langle k+q|\rho_{\mathbf{n}_{qm}^{-}}^{(n-1,\mathrm{eq})}|k+k_{\mathbf{n}}\rangle.\end{split}$ (75)
In contrast to the HEOM for single-particle properties [Eq. (62)], the closing term in Eq. (68) is crucial in avoiding the long-time numerical instabilities characteristic for electron–phonon models with a finite number of phonon modes [49, 50, 60]. We always choose a sufficiently large $D$ so that the closing term merely stabilizes the long-time dynamics of the zeroth- and first-tier auxiliaries entering Eqs. (70)–(72) without compromising the numerical exactness of the HEOM mobility [50]. We have also checked that using other closing terms [61, 59] does not change the mobility obtained by propagating Eq. (68[50, 51].
@@ IV.1.1 Quasiparticle properties
The results are shown in Fig. 4. Since most points in Fig. 4 lie in the higher-temperature range, where the mobility $\mu$ is small, we chose to plot $1/\mu(T)$ for better clarity. For completeness, the plot of $\mu(T)$ is provided in the following subsection; see Fig. 6($a_{2}$)–($c_{2}$). The curves labeled ”HEOM full” and ”HEOM IPA” represent numerically exact benchmarks for the mobility with [Eqs. (67)–(75)] and without [Eqs. (65) and (66)] vertex corrections, respectively. As seen from the figure, the discrepancy between the two clearly demonstrates that vertex corrections are not small, although the IPA result remains of the same order of magnitude as the full result. This difference is not solely due to the fact that $\mu_{x}$ vanishes within IPA [see Eq. (11)]; it is also evident in both the electronic [Panel (b)] and phononic [Panel (c)] contributions. In particular, while the IPA captures the qualitative temperature dependence of the electronic term in Panel (b), the phonon-assisted contribution in Panel (c) is not even qualitatively reproduced. Equation (66) reveals that the overall temperature dependence of $\mu_{ph,\mathrm{HEOM}}^{\mathrm{IPA}}$ is determined by the competition of the temperature effects on the phonon propagator and on the product of particle and hole propagators [see also Fig. 1(b)]. While the former promote $\mu_{ph,\mathrm{HEOM}}^{\mathrm{IPA}}$ increasing with temperature due to the enhanced phonon occupation, and dominate at lower temperatures, the latter promote $\mu_{ph,\mathrm{HEOM}}^{\mathrm{IPA}}$ decreasing with temperature, similarly to $\mu_{e,\mathrm{HEOM}}^{\mathrm{IPA}}$ [see Eq. (65) and Fig. 4(b)], and dominate at higher temperatures. Therefore, $\mu_{ph,\mathrm{HEOM}}^{\mathrm{IPA}}$ reaches its maximum in the intermediate-temperature range. In addition to the above-discussed Wick-decoupled contribution [see also Eq. (12)], the full solution $\mu_{ph,\mathrm{HEOM}}^{\mathrm{full}}$ contains both the genuine particle–hole correlations, which are behind the difference between $\mu_{e,\mathrm{HEOM}}^{\mathrm{full}}$ and $\mu_{e,\mathrm{HEOM}}^{\mathrm{IPA}}$ (and also the vertex corrections in the Holstein model [46]), and genuine electron–phonon correlations, which are formally embodied in the first- and second-tier HEOM auxiliaries [50]. These all stem from the cluster expansion-based decomposition [67] of the two-electron–two-phonon correlation function
$\langle c^{\dagger}_{\bf k_{2}+q_{2}}(t)c_{\bf k_{2}}(t)X_{\bf q_{2}}(t)c_{\bf k_{1}+q_{1}}^{\dagger}c_{\bf k_{1}}X_{\bf q_{1}}\rangle$
determining $\mu_{ph}$. In Sec. IV.1.3, we argue that the electron–phonon correlations are ultimately responsible for turning the nonmonotonic temperature dependence of $\mu_{ph,\mathrm{HEOM}}^{\mathrm{IPA}}$ into the monotonic decrease of $\mu_{ph,\mathrm{HEOM}}^{\mathrm{full}}$ with temperature. We finally note that, for similar values of $\omega_{0},\lambda,$ and $T$, vertex corrections were found to be less pronounced in the Holstein model [46]. In contrast, their stronger impact in the Peierls model can be attributed to the pronounced momentum dependence of the electron–phonon matrix elements $g_{k,q}$ in both $k$ and $q$ [see Eq. (25)], as well as to the genuine electron–phonon correlations, which are not present in models without the phonon-assisted current.
@@
All other methods (MA, SCMA, CE) are calculated within IPA, and thus need to be compared with HEOM IPA. The above rationalization of the temperature dependence of $\mu_{ph}$ within the IPA is general and not limited to the HEOM method. Indeed, Fig. 4(c) reveals that MA, SCMA, and CE capture the nonmonotonic temperature dependence of the HEOM IPA benchmark. However, the MA shows quantitative discrepancies when compared to HEOM IPA, while the functional form of $\mu(T)$ at high temperatures, where the phonon-assisted contribution dominates, predicts a significantly steeper increase of $1/\mu$ with temperature than the benchmark. To improve upon the predictions of MA, one can use either SCMA or CE. Both of them are in very good agreement with HEOM IPA. However, while SCMA has self-consistency, CE is completely one-shot. This property makes it a computationally efficient and appealing alternative to other methods.
@@ IV.1.3 Mobility results in a wide range of parameter regimes
In Figs. 6 and 7, we present the mobility results over a wide range of parameter regimes. We do not discuss each panel individually, as the main conclusions are analogous to those from the previous subsection. Specifically: i) Vertex corrections are clearly visible but do not alter the order of magnitude of the results. In contrast to the Holstein model [46], however, we now also observe non-negligible vertex corrections even at weak interaction strengths [see, for example, Fig. 7($a_{1}$)]. This can be attributed to the strong dependence of the matrix elements $g_{k,q}$ on both $k$ and $q$. Additionally, in contrast to the intermediate- and strong-interaction regimes, the IPA qualitatively captures the temperature dependence of $\mu_{ph}$ in the weak-interaction regime [see, for example, Fig. 6($c_{1}$)]. Then, one expects that the genuine electron–phonon correlations are weak, so that the quantitative discrepancies between the HEOM full and HEOM IPA results for both $\mu_{ph}$ and $\mu_{e}$ originate from the particle–hole correlations that the IPA misses. Their absence does not alter the correct temperature dependence of $\mu_{e}$, at least in the ranges of the temperature and interaction strength we examine [see Figs. 6($b_{1}-b_{4}$) and 7($b_{1}-b_{4}$)], and the same then applies to $\mu_{ph}$ [see Fig. 6($c_{1}$)]. The above reasoning then strongly suggests that the qualitatively incorrect temperature dependence of $\mu_{ph}$ within the IPA [see Figs. 6($c_{2}-c_{4}$) and 7($c_{2}-c_{4}$)] is due to the neglect of genuine electron–phonon correlations, and not due to the neglect of particle–hole correlations. ii) The CE approach successfully reproduces the correct IPA result well beyond the range of validity of the Boltzmann quasiparticle picture. Unlike CE, which is a one-shot method, SCMA incorporates self-consistency, allowing it to achieve accurate results that also agree with the numerically exact HEOM IPA benchmark. Overall, both CE and SCMA provide a clear improvement over MA, which often deviates from the exact result except in the very weak coupling regime.
At lower temperatures, it becomes increasingly difficult to converge the CE results (this issue does not arise within SCMA). We verified that this problem is related to the long tail of the CE spectral functions extending toward negative frequencies, in close analogy with Fig. 5(c); see the corresponding spectral functions in Sec. B of SM [52]. We also observed that this issue is even more pronounced at stronger coupling strengths. This behavior is expected, as the spectral functions become broader in this regime, thereby enhancing the contribution of the tail.
In these parameter regimes, the following issue arises. Although it is sometimes possible, with considerable effort, to converge the CE mobility results with respect to the frequency cutoff $\Lambda$ at stronger couplings and lower temperatures, CE typically underestimates the HEOM IPA benchmark in these cases [see, for example, Fig. 6($a_{4}$)]. This can also be understood as a direct consequence of the spectral function tail. As discussed at the end of Sec. IV.1.2, the tail has the largest impact on $n_{e}$ (see Eq. (4)), making it unphysically large. Since $n_{e}$ appears in the denominator of the expressions for $\mu_{e}$ and $\mu_{ph}$ [see Eqs. (10) and (13)], these quantities within CE are systematically underestimated compared to the exact solution.